To burn 456 grams of nonane (C9H20), how many liters of oxygen are needed?


1
Expert's answer
2021-02-04T07:07:12-0500

C9H20+14O2=9CO2+10H20

"n(C9H20)=\frac{m}{M}=\frac{456}{128.259}=3.56"

"\frac{3.56}{1}=\frac{x}{14}"

x=49.77

"V(O2)=Vm\times n=22.4\times49.77=1114.94" L